The value of x + x(xx) when x = 2 is:

(a) 10, (b) 16, (c) 18, (d) 36, (e) 64

 

Grade:11

3 Answers

Samyak
7 years ago
Hello 
 
Your question answer can be given as 
 
x+x(xx) = 2+ 2(22)
=10
Please feel free to post any doubts you have. This forum is for clearing the doubts only
sanjoy
7 years ago
it is class 7 level qsn...thle answer will be 10...ask more difficult qsns...ok                                                                                                                                                                                                   
2+2(2​2​)=10...or any doubts..thn ask....
 
Abhishek kumar
6 years ago
According to your questionThe value of x=2 So, x+x(x x) x+x(x^2)Now, 2+2(2^2) =2+2(4) =2+8 =10 So, the answer is 10